Show that the sequence (fn) sequence where


fn(x)= x/(1+nx^2), x∈[2,∞] is uniformly convergent in [2,∞]

1
Expert's answer
2022-01-30T13:15:37-0500

To prove that the given sequence of functions converges uniformly on "\displaystyle
[2,\infty)", it suffices to show that the sequence of functions converges uniformly on "\displaystyle \R".

Now,

"\displaystyle
\left|f_n(x)\right|=\left|\frac{x}{1+nx^2}\right|=\left|\frac{x\sqrt{n}}{(1+nx^2)\sqrt{n}}\right|=\frac{1}{\sqrt{n}}\left(\frac{\sqrt{n}\ |x|}{1+{(\sqrt{n}|x|)}^2}\right)=\frac{1}{\sqrt{n}}\left(\frac{t}{1+t^2}\right)"

where "\displaystyle
t=\sqrt{n}\ |x|.\text{ But, }\frac{t}{1+t^2}\leq\frac{1}{2}\ \ \ \forall\ t\in\R\ \text{since } (1−t)^2 ≥ 0,\ \Rightarrow 2t ≤ 1 + t^2."

Thus, from the above inequality, we have that; "\displaystyle
|f_n(x)|=\frac{1}{\sqrt{n}}\left(\frac{t}{1+t^2}\right)≤\frac{1}{\sqrt{n}}\left(\frac{1}{2}\right)\rightarrow0,\ \text{as } n\rightarrow \infty\ \ \ \forall\ \ \ x\in\R.\\
\text{Hence, given }\epsilon>0\ \text{choose } N = \frac{1}{4\epsilon^2}.\ \text{Then } |f_n(x)| < \epsilon\ \ \ \forall\ \ x\in\R\ \text{if}\ n>N." Showing that the sequence of functions converges uniformly to "\displaystyle
0 \text{ on }\R." And since "\displaystyle
[2,\infty)\subset\R," then the given sequence of functions converges uniformly to "\displaystyle
0 \text{ on }[2,\infty)."